OpenStudy (anonymous):

Someone PLEASE help me i've been stuck on this for days. 2. A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 1. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. 2. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). 3. Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

OpenStudy (anonymous):

@ganeshie8 @ash2326 someone?

OpenStudy (ash2326):

Let's solve this, can you multiply the first equation by any integer other than 1 ?

OpenStudy (anonymous):

yes, I did by 5 and got 10x+35y=5

OpenStudy (ash2326):

Good, can you write the modified first equation and the original second equation?

OpenStudy (anonymous):

yes 10x+35y=5 and -3x-4y=5

OpenStudy (ash2326):

Ok, do you know any method of solving equations with 2 variables?

OpenStudy (anonymous):

yeah

OpenStudy (ash2326):

can you try to use it to solve these equations?

OpenStudy (anonymous):

I'm just confused about what they're trying to tell me and I have troubles adding those type of equations together, I asked my teacher but it wasn't much help.

OpenStudy (ash2326):

10x+35y=5 and -3x-4y=5 ok, multiply the first equation by 3 and second by 10, can you try?

OpenStudy (anonymous):

yes one second

OpenStudy (ash2326):

ok

OpenStudy (anonymous):

for the first one it would be 30x+105y=15 and the second one it would be -30x-40y=50 I believe...

OpenStudy (anonymous):

am i correct?

OpenStudy (ash2326):

yes let's write it as 30x+105y=15 -30x-40y=50 Can you add both the equations?

OpenStudy (anonymous):

I don't think so :/

OpenStudy (ash2326):

why? just add left side together and right side together, simple addition :)

OpenStudy (anonymous):

i'll try

OpenStudy (ash2326):

good

OpenStudy (anonymous):

umm I got 0+65y=65

OpenStudy (ash2326):

yes, good, can you divide both sides by 65?

OpenStudy (anonymous):

yes it would be 1

OpenStudy (anonymous):

so y=1?

OpenStudy (anonymous):

are you still there

OpenStudy (ash2326):

sorry I was afk, yes good. Can you put y in any of the equations and find the value of y?

OpenStudy (anonymous):

in which equations

OpenStudy (ash2326):

any of the two equations, the modified first or the original second

OpenStudy (anonymous):

well I don't believe so...

OpenStudy (ash2326):

try, put y=1 and solve for x

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

for 30x+105y=15 I got x= -3

OpenStudy (ash2326):

yes, correct. So we are done with 2 parts

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

so for the first part I write the modified question and also write what the answer would be when i add them?

OpenStudy (ash2326):

first one the modified and the original second equation second part solving for x and y for the third part solve the original equations to find x and y, they would come the same and you'll prove that

OpenStudy (anonymous):

okay just to be sure... to solve the original equations i have to add them together first right?

OpenStudy (ash2326):

no need, just multiply both of them with integers such that coefficient of x or y in both the equations is same. 2x + 7y = 1 -3x – 4y = 5

OpenStudy (anonymous):

that kinda confused me

OpenStudy (ash2326):

coefficents of x 2 and -3 you can multiply first by 3 and second by 2, you'll get 6x+21y=3 -6x-8y=10 you can solve by adding which will eliminate x and then you can find y or you can multiply first by 4 and second by 7 and add them to eliminate y anything would work

OpenStudy (anonymous):

okayy, thank you soo much! :) i can catch back up on pace now lol

OpenStudy (ash2326):

yeah :)